OpenStudy (anonymous):

If a = 3 3 in the right triangle below, what is the value of b?

OpenStudy (anonymous):

it's number 84

OpenStudy (paxpolaris):

what do you think it is...

OpenStudy (anonymous):

B

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no its A = 9

OpenStudy (paxpolaris):

how do you get that?

OpenStudy (anonymous):

me? or Nira

OpenStudy (paxpolaris):

@Hm1581

OpenStudy (paxpolaris):

you must have used the ratios....

OpenStudy (anonymous):

yeah..i think i might have..

OpenStudy (anonymous):

first thing u notice is its a 30-60-90 triangle then u kno that side opposite to 30 is 1/2 hypotenuse (c) so 1/2 (c) = 3 rad 3 so c = 6 rad 3 then u kno that side opp to 60 is (rad 3) /2 so b = rad 3 /2 * 6 rad 3 = 9

OpenStudy (anonymous):

Ok! thank soo much nira!!!! :)

OpenStudy (paxpolaris):

\(6sqrt3\) is the value of the hypotenuse

OpenStudy (anonymous):

ok, thanks pax :)